Which of the following will provide a better yield. (Interest rates are annual.)
Plan A: $40,000 invested for 3 years at 4.5% compounded quarterly
Plan B: $40,000 invested for 3 years at 4.4% compounded continuously


Answer by rfer(10417)

A=P(1+r/n)^t*n
A=40000(1+.045/4)^3*4
A=40000(1.01125)^12
A=40000(1.1436774)
A=$45,747.10
-----------------
A=P(e)^rt
A=40000(2.7183)^.044*3
A=40000(2.7183)^.132
A=40000(1.1411)
A=$45,644.37